OpenStudy (anonymous):

FAN FOR ANSWER! Which of the following exponential functions goes through the points (1, 12) and (2, 36)? f(x) = 4(3)−x f(x) = 3(4)−x f(x) = 3(4)x f(x) = 4(3)x

OpenStudy (welshfella):

plug in f(x) = 12 and x = 1 and see if they fit do the same for f(x) = 36 and x = 2

OpenStudy (anonymous):

Not so sure how to do that give me one second...

OpenStudy (anonymous):

Oh I see let me figure it out really quickly...

OpenStudy (anonymous):

Wait never mind I have no idea lmao...

OpenStudy (welshfella):

I'm a little confused here which numbers/letters are the exponents?

OpenStudy (anonymous):

Hold on I gave it to you in a weird way.

OpenStudy (welshfella):

4(3) - x = 4^3 - x or 4^(3-x) ?

OpenStudy (anonymous):

Which of the following exponential functions goes through the points (1, 12) and (2,36)? f(x) = 4(3)−x f(x) = 3(4)−x f(x) = 3(4)x f(x) = 4(3)x

OpenStudy (anonymous):

So... I think your right...

OpenStudy (welshfella):

which one is it?

OpenStudy (anonymous):

To be honest Im really confused rn Im sorry...

OpenStudy (welshfella):

3(4)x could mean 3^4x or 3(4)^x !!!

OpenStudy (anonymous):

Oh ok but which one is it! lol how about another question?

OpenStudy (welshfella):

3(4)^x would it the point (1,12)

OpenStudy (welshfella):

* fit

OpenStudy (welshfella):

= but not (2,36)

OpenStudy (anonymous):

Ok i'm not trying to piss you off but do you mind elaborating?

OpenStudy (anonymous):

For example I'm not the greatest at math so how would I come across that answer?

OpenStudy (welshfella):

no i'm too tired to be honest..

OpenStudy (anonymous):

Its fine lol go take a break!